x/16 + 3x/8 = x/12
Answer by stanbon(75887)

x/16 + 3x/8 = x/12
------------
Multiply thru by 48 to get:
3x + 18x = 4x
17x = 0
x = 0
===========================
Cheers,
Stan H.